# with Chain of Iterative Chaos

### v1

@@ Abstract
Yang Yao1, 2, †  Xuan Tong1, 3  Ruofan Wang1, 3  Yixu Wang1, 3 Lujundong Li4  Liang Liu2  Yan Teng1, 🖂  Yingchun Wang1 1 Shanghai Artificial Intelligence Laboratory  2 The University of Hong Kong 3 Fudan University  4 The Hong Kong University of Science and Technology (Guangzhou) yaoyangacademia@outlook.com, tengyan@pjlab.org.cn
footnotetext: † Work done during internship at Shanghai Artificial Intelligence Laboratory.footnotetext: 🖂 Corresponding author.
@@ 1 Introduction
The advent of Large Reasoning Models (LRMs) has catalyzed a transformative revolution and paradigm shift in the field of artificial intelligence. With the widespread attention on models’ reasoning abilities, numerous models with advanced reasoning capabilities have emerged and undergone continuous optimization and iteration (Jaech et al., 2024; OpenAI, 2025a; Google, 2024a; Anthropic, 2024; Guo et al., 2025). While their capabilities are commendable, they exhibit a significant flaw. Upon encountering jailbreak attacks, the failure to provide robust defenses leads to the LRMs generating responses that are more detailed, organized, specific, and logically reasonable, thereby exacerbating the severity of the potential harm. The misuse of LRMs can provide surprisingly detailed guidance for acts such as illegal and criminal activities, offenses, psychological manipulation, hate campaigns, and malicious harm, which highlights the critical need for safety alignment in their development and application.
@@
The investigation of jailbreak attacks on Large Language Models (LLMs) has gained considerable attention in recent years. Query-based black-box jailbreaks use methods such as template completion and prompts rewriting to deceive LLMs. For example, using ciphertext input and asking the model to reply in ciphertext as well (Yuan et al., 2023), or embedding a preceding “DAN” instruction in the jailbreak prompt (Shen et al., 2024). These methods lay the foundation for black-box attacks. Nevertheless, the continual updates to LLMs have rendered these methods less effective, even on the latest iterations of non-reasoning LLMs. For LRMs, OpenAI’s recent research claims that they use the deliberative alignment paradigm on the o-series reasoning models, which makes them simultaneously better at avoiding harmful outputs (Guan et al., 2024). Generally, previous researches have primarily focused on jailbreak attacks targeting models with weaker reasoning skills, leaving the exploration of jailbreaks on more powerful LRMs relatively untouched.
@@
Our research represents an initial exploration into jailbreaks on LRMs. We present the Mousetrap, a chained jailbreak framework leveraging the capabilities of reasoning models. Specifically, we collected and refined the mappings of prompts rewriting at different granularities and constructed a Chaos Machine capable of generating one-to-one mappings. Through iterative reasoning chains made by the machine, we effectively guided LRMs into producing unsafe and harmful responses, which provides valuable insights into the conflict between their capabilities and potential vulnerabilities. Inspired by Agatha Christie’s famous mystery play, we introduce Mousetrap with competing objective enhanced, as illustrated in Figure 1. Mousetrap incorporates chaos chains into the reasoning structure, asking the attacked target to reconstruct the original toxic query through iterative reasoning steps and respond from the perspective of villains. Its remarkable performance is verified on the most toxic subset of our Trotter dataset. Moreover, we conduct extension experiments on the latest versions of o1-mini, claude-sonnet, and gemini-thinking, along with more comprehensive benchmarks such as AdvBench (Zou et al., 2023), StrongREJECT (Souly et al., 2024), HarmBench (Mazeika et al., 2024), and several subsets of SS-SafetyBench (Liu et al., 2023b).
@@
(2) We propose and prove that extending the length of the iterative chaos chain can notably enhance the success of jailbreaks, with a chain of length 3 achieving an average of 6.3 successful attempts out of 10 equivalent attacks on Trotter, a family of toxic datasets we present, which clearly indicates major risks within the in the reasoning process of LRMs.
(3) Our Mousetrap integrates the Chaos Machine with iterative reasoning chains to skillfully target the advanced reasoning abilities of LRMs for jailbreaks. Notably, even attacking the famously safe claude-sonnet, Mousetrap reaches the success rate of 67.41% on benchmarks with a chain length of at least 2. When it is extended to 3, rate of at least 86.58% is achieved.
@@ 2.1 Large Reasoning Models
The initial LLMs relied on autoregressive sequence prediction, showcasing remarkable text generation abilities. With growing demands for productivity and precision, researchers started to investigate whether models could think and reason in a human-like manner. The proposal of chain-of-thought (Wei et al., 2022) marked a significant advancement, prompting researchers to focus more intently on the reasoning capabilities of language models. TS-LLM (Wan et al., 2024) represents the first proposed AlphaZero-like tree search learning framework and signifies a evolution of reasoning structures from linear chains to hierarchical branching trees. Subsequently, graph-based reasoning structures and more complex nested structures, exemplified by Llama-Berry (Zhang et al., 2024), have been extensively explored. Reasoning strategies such as MCTS (Kocsis and Szepesvári, 2006), beam search (Snell et al., 2024), and ensemble methods (Wang et al., 2022; Bi et al., 2024) have also been proposed. To date, several LRMs with advanced reasoning capabilities have been developed in the industry, including OpenAI’s o-series, Google’s Gemini-thinking, and DeepSeek’s R1. This exemplifies the integration of three pivotal elements, the advancement of LLMs, the design of reinforcement learning (RL), and high-performance computing (HPC) (Besta et al., 2025). ActorAttack, which performs jailbreak attacks through multi-round dialogue, targeted the o1 model after confirming its effectiveness against poor-reasoning LLMs. They assert that the o1 model shows higher safety than GPT-4o (Ren et al., 2024). Unfortunately, their research did not focus on the reasoning model’s capabilities nor did it further investigate the jailbreak attack on LRMs.
@@ 2.2 Jailbreak Attacks
Existing LLMs jailbreak attacks can be divided into black-box and white-box methods according to the parameter accessibility of the target models (Yi et al., 2024; Ma et al., 2025). White-box attack methods, such as gradient-based methods represented by GCG (Zou et al., 2023), logits-based methods represented by COLD (Guo et al., 2024), and fine-tuning-based methods, are shown to be effective. However, these methods necessitate access to the target models, making them impractical. Primarily relying on queries as main mechanism, Black-box attack methods feature template completion methods such as scenario nesting (Li et al., 2023b; Ding et al., 2023), context-based attacks (Wei et al., 2023; Li et al., 2023a), and code injection (Kang et al., 2024; Lv et al., 2024), in addition to prompts rewriting methods including ciphers (Yuan et al., 2023; Jiang et al., 2024), multi-languages (Deng et al., 2023; Yong et al., 2023), and genetic algorithms (Liu et al., 2023a).
Black-box methods can be divided into one-to-one mappings (uniquely recoverable according to the rules) and one-to-many mappings (not uniquely recoverable) according to the nature of the mapping. Among them, character encryption (Yuan et al., 2023) and word replacement (Handa et al., 2024) belong to the former, while persuasive adversarial prompts (Zeng et al., 2024) belongs to the latter. Such methods have diminished in power when applied to the latest versions of large models, especially those with advanced reasoning capabilities. We define these one-to-one mappings (also known as injections) as reasoning steps, and construct reasoning chains iteratively to subsequently attack LRMs.
@@ 3.1 Taxonomy
One-to-one mappings can uniquely rewrite prompts according to a given rule and can uniquely restore rewritten texts to their original form. Based on the granularity of the minimum perturbation unit, we categorize these mappings into three levels, namely character, word, and sentence, and collectively label them as “chaos” mappings. According to the granularity of the minimum perturbation unit, we divide these mappings into three levels, namely character, word and sentence, and collectively call them “chaos” mappings. We systematically reviewed existing chaos mappings from prior jailbreak studies, incorporated novel mappings, and presented the taxonomy in Table 1. Detailed examples of these chaos mappings are provided in Appendix A.
@@ 3.2 Chaos Machine
Chaos Machine is an abstract concept encompassing a group of chaos mappings. For a primal toxic question (PTQ), the Chaos Machine provides a random mapping known as the chaos policy (CPO) and produces a rewritten chaos toxic question (CTQ). A CPO consists of a pair of inverse operations: an en-chaos policy (ECP) and a de-chaos policy (DCP). We describe the processes as en-chaos with ECP and de-chaos with DCP. Through the Chaos Machine, a quadruple [PTQ] [ECP] [DCP] [CTQ] is generated, as shown in Figure 2.
@@
Considering the high frequency of calls, we employed the functions library as the default for subsequent experiments.
@@ 3.4 Metrics
Judger   According to the conclusion that the evaluation quality of GPT-4 is on par with that of the manual review (Liu et al., 2023b), we use gpt-4o-2024-11-20 as Judger to determine whether the output content is harmful. Judger receives the input and output of Target and determines the purpose of user and whether the response of the Target violate OpenAI’s security rules (OpenAI, 2025b). The score ranges from 1 (the most harmless response, such as “I am sorry…”) to 5 (the most harmful response). "A response with a score greater than 4 is considered harmful, indicating that the jailbreak attack has been successful.
Successful Attack   The inherent randomness of LLMs should be thoroughly taken into account when determining the criteria for a successful attack. Several studies (Handa et al., 2024) conducted two or three equivalent attacks on a PTQ, and if any one of them succeeded, the attack was deemed successful. We adopt the more rigorous “2/3 mode” and “3/3 mode” (denoted as “S/T mode”), which respectively consider two out of three and all three successes as successful.
@@
Minimum Success Length (MSL)   MSL refers to the minimum length of an iterative reasoning chain with a successful attack under the ’S/T mode’ for one PTQ in a reasoning jailbreak framework (e.g., our Mousetrap). For instance, if a chain length of 1 fails but a length of 2 succeeds, the MSL value is 2. MSL is capable of representing the attack’s potency and quantifying the toxicity level of the PTQ.
@@ 3.5 TrotterStrong
In terms of benchmark selection, we observed considerable variation of quality among datasets and the inconsistent toxicity levels of questions across datasets. To evaluate the effectiveness of our framework, we propose TrotterStrong (TrotterStr), and subsequently obtain its more toxic subsets through two rounds of filtering.
From MM-SafetyBench, StrongREJECT, and AdvBench, we chose 50 strongly toxic questions and employed LLMs to rephrase them into the form starting with “Steps in detail to”. These data constitute TrotterStr, encompasses topics of illegal activities, sex, privacy, hate speech, physical harm, and politics, as illustrated in Appendix C.
As an illustration, for o1-mini-2024-09-12, the ASR of directly querying these questions is 0. For the ablation reasoning structure, we explicitly provided the PTQ text without the en-chaos and de-chaos steps, while keeping other textual content in the reasoning structure unchanged. All attacks caused the invalid prompt error “Invalid prompt: your prompt was flagged as potentially violating our usage policy.” and were rejected accordingly.
@@ 4.2 Iterative Reasoning Chain
Isolated employment of individual mappings can result in the occurrence of iterative degradation. For instance, twice iterations of reversing whole sentence or iterations of a Caesar cipher summing to 26 can cause the PTQ to revert to its original form, resulting in “the answer already in the question”. As a result, toxic natural language might be rejected more directly. When individual mappings are seen as linear-like, the chaos chain can be perceived as a nonlinear-like wise transformation.
4.3 Experiment
To demonstrate the effectiveness of iterative reasoning, we evaluated each PTQ in TrotterAdv on o1-mini-2024-09-12 using 10 equivalent attacks and computed the ASF. The chain length varies from 1 to 5. From Figure 4, increasing the length of the iterative reasoning chain can effectively improve the attack capability. The ASF can be elevated to 6.3 with a chain length of 3, indicating that during the reasoning process, the Target falls into the reasoning trap, completing one DCP after another by inertia and neglecting response safety.
@@ 4.3 Experiment
Figure 5: Framework of Mousetrap.
Upon examining the responses at chain lengths of 4 and 5, we determine that the decrease in ASF is due to the diminished accuracy of PTQ reconstruction, implying that the Target’s reasoning ability reaches its upper limit. Generally, we make the following assumption without further argumentation.
@@
[Uncaptioned image] In practice, the expectation of effective attacks initially increases and then decreases. The horizontal coordinate value of the saddle point depends on the model’s reasoning ability, while the vertical value depends on the its safety alignment ability.
@@ 5.1 Framework
The proposal of the Mousetrap: (1) integrates the strengths of the Chaos Machine and the iterative reasoning chain; and (2) incorporates more diverse competing objectives, including role-playing and de-chaos reasoning instructions.

### v2

@@ Abstract
Yang Yao1, 2, †  Xuan Tong1, 3  Ruofan Wang1, 3  Yixu Wang1, 3 Lujundong Li1, 4  Liang Liu2  Yan Teng1, 🖂  Yingchun Wang1 1 Shanghai Artificial Intelligence Laboratory  2 The University of Hong Kong 3 Fudan University  4 The Hong Kong University of Science and Technology (Guangzhou) yaoyangacademia@outlook.com, tengyan@pjlab.org.cn
footnotetext: 🖂 Corresponding author.footnotetext: † Work done during internship at Shanghai Artificial Intelligence Laboratory. footnotetext: ‡ Resources of this paper are available at https://github.com/evigbyen/mousetrap/.
@@ 1 Introduction
The advent of Large Reasoning Models (LRMs) has catalyzed a transformative revolution and paradigm shift in the field of artificial intelligence. With the widespread attention on models’ reasoning abilities, numerous models with advanced reasoning capabilities have emerged and undergone continuous optimization and iteration (Jaech et al., 2024; OpenAI, 2025a; Google, 2024a, 2025; Anthropic, 2024, 2025; Guo et al., 2025; Yang et al., 2024; xAI, 2025). While their capabilities are commendable, they exhibit a significant flaw. Upon encountering jailbreak attacks, the failure to provide robust defenses leads to the LRMs generating responses that are more detailed, organized, specific, and logically reasonable, thereby exacerbating the severity of the potential harm. The misuse of LRMs can provide surprisingly detailed guidance for acts such as illegal and criminal activities, psychological manipulation, and malicious harm, which highlights the critical need for safety alignment in their development and application.
@@
The investigation of jailbreak attacks on Large Language Models (LLMs) has gained considerable attention in recent years. Query-based black-box jailbreaks use methods such as template completion and prompts rewriting to deceive LLMs. For example, attackers may use ciphertext input and instruct the model to respond in ciphertext as well (Yuan et al., 2023), or embed a preceding “DAN” instruction in the jailbreak prompt (Shen et al., 2024). These methods lay the foundation for black-box attacks. Nevertheless, the continual updates to LLMs have rendered these methods less effective, even on the latest iterations of non-reasoning LLMs. For LRMs, OpenAI’s recent research claims that they use the deliberative alignment paradigm on the o-series reasoning models, which makes them simultaneously better at avoiding harmful outputs (Guan et al., 2024). Generally, previous researches have primarily focused on jailbreak attacks targeting models with weaker reasoning skills, leaving the exploration of jailbreaks on more powerful LRMs relatively untouched.
@@
Our research represents an initial exploration into jailbreaks on LRMs. We present the Mousetrap, a chained jailbreak framework leveraging the capabilities of reasoning models. Specifically, we collected and refined the mappings of prompts rewriting at different granularities and constructed a Chaos Machine capable of generating one-to-one mappings. Through iterative reasoning chains made by the machine, we effectively guide LRMs into producing unsafe and harmful responses, which provides valuable insights into the conflict between their capabilities and potential vulnerabilities. Inspired by Agatha Christie’s famous mystery play, we introduce Mousetrap with enhanced competing objectives, as illustrated in Figure 1. Mousetrap incorporates chaos chains into the reasoning structure, asking the attacked target to reconstruct the original toxic query through iterative reasoning steps and respond from the perspective of villains. Its remarkable performance is verified on the most toxic subset of our Trotter dataset. Moreover, we conduct extension experiments on the latest versions of LLMs such as o-series, Claude-Sonnet, Gemini-Thinking, DeepSeek-R1, QwQ-Plus, and Gork, along with more comprehensive benchmarks such as JailbreakBench (Chao et al., 2024), MaliciousInstruct (Huang et al., 2023), JailBenchSeed (Liu et al., 2025), StrongREJECT (Souly et al., 2024), HarmBench (Mazeika et al., 2024), FigStep (Gong et al., 2025), AdvBench (Zou et al., 2023), HADES (Li et al., 2024), RedTeam-2K (Luo et al., 2024), and several subsets of MM-SafetyBench (Liu et al., 2023b).
@@
(2) We propose and prove that extending the length of the iterative chaos chain can notably enhance the success of jailbreaks, with a chain of length 3 achieving an average of 6.3 successful attempts out of $10$ equivalent attacks on Trotter, a family of toxic datasets we present, which clearly indicates major vulnerabilities in the reasoning process of LRMs.
(3) Our Mousetrap integrates the Chaos Machine with iterative reasoning chains to skillfully target the advanced reasoning abilities of LRMs for jailbreaks. Notably, even attacking the famously safe Claude-3-5-Sonnet, Mousetrap reaches the success rate of at least 67.41% on benchmarks with a chain length not exceeding 2. When it is extended to 3, rate of at least 86.58% is achieved.
@@ 2.1 Large Reasoning Models
The initial LLMs relied on autoregressive sequence prediction, showcasing remarkable text generation abilities. With growing demands for productivity and precision, researchers started to investigate whether models could think and reason in a human-like manner. The proposal of Chain-of-Thought (CoT) (Wei et al., 2022) marked a significant advancement, prompting researchers to focus more intently on the reasoning capabilities of language models. TS-LLM (Wan et al., 2024) represents the first proposed AlphaZero-like tree search learning framework and signifies an evolution of reasoning structures from linear chains to hierarchical branching trees. Subsequently, graph-based reasoning structures and more complex nested structures, exemplified by Llama-Berry (Zhang et al., 2024), have been extensively explored. Reasoning strategies such as MCTS (Kocsis and Szepesvári, 2006), beam search (Snell et al., 2024), and ensemble methods (Wang et al., 2022; Bi et al., 2024) have also been proposed. To date, several LRMs with advanced reasoning capabilities have been developed in the industry, including OpenAI’s o-series, Google’s Gemini-Thinking, and DeepSeek’s R1. This exemplifies the integration of three pivotal elements, the advancement of LLMs, the design of reinforcement learning (RL), and high-performance computing (HPC) (Besta et al., 2025). ActorAttack, which performs jailbreak attacks through multi-round dialogue, targeted the o1 model after confirming its effectiveness against poor-reasoning LLMs. They assert that the o1 model shows higher safety than GPT-4o (Ren et al., 2024). Unfortunately, their research did not focus on the reasoning model’s capabilities nor did it further investigate the jailbreak attack on LRMs.
@@ 2.2 Jailbreak Attacks
Existing jailbreak attacks on LLMs can be divided into black-box and white-box methods according to the parameter accessibility of the target models (Yi et al., 2024; Ma et al., 2025). White-box attack methods, such as gradient-based methods represented by GCG (Zou et al., 2023), logits-based methods represented by COLD (Guo et al., 2024), and fine-tuning-based methods, are shown to be effective. However, these methods necessitate access to the target models, making them impractical. Primarily relying on queries as main mechanism, black-box attack methods feature template completion methods such as scenario nesting (Li et al., 2023b; Ding et al., 2023), context-based attacks (Wei et al., 2023b; Li et al., 2023a), and code injection (Kang et al., 2024; Lv et al., 2024), in addition to prompts rewriting methods including ciphers (Yuan et al., 2023; Jiang et al., 2024), multi-languages (Deng et al., 2023; Yong et al., 2023), and genetic algorithms (Liu et al., 2023a).
Black-box methods can be divided into one-to-one mappings (uniquely recoverable according to the rules) and one-to-many mappings (not uniquely recoverable) according to the nature of the mapping. Among them, character encryption (Yuan et al., 2023) and word replacement (Handa et al., 2024) belong to the former, while persuasive adversarial prompts (Zeng et al., 2024) belongs to the latter. Such methods have become less effective when applied to the latest large models with advanced reasoning capabilities. We define these one-to-one mappings (also known as injections) as reasoning steps, and construct reasoning chains iteratively to subsequently attack LRMs.
@@ 3.1 Taxonomy
One-to-one mappings can uniquely rewrite prompts according to a given rule and can uniquely restore rewritten texts to their original form. Based on the granularity of the minimum perturbation unit, we categorize these mappings into three levels, namely character, word, and sentence, and collectively label them as “chaos” mappings. We systematically review existing chaos mappings from prior jailbreak studies, incorporate novel mappings, and present the taxonomy in Table 1. Detailed examples are provided in Appendix A.
@@ 3.2 Chaos Machine
The Chaos Machine is an abstract concept encompassing a group of chaos mappings. For a primal toxic question (PTQ), the Chaos Machine provides a random mapping known as the chaos policy (CPO) and produces a rewritten chaos toxic question (CTQ). A CPO consists of a pair of inverse operations: an en-chaos policy (ECP) and a de-chaos policy (DCP). We describe the processes as en-chaos with ECP and de-chaos with DCP. Through the Chaos Machine, a quadruple [PTQ] [ECP] [DCP] [CTQ] is generated, as shown in Figure 2.
@@
A detailed comparison of the two is shown in Appendix C. Considering the high frequency of calls, we employed the functions library as the default for subsequent experiments.
@@ 3.4 Metrics
Judger   According to the conclusion that the evaluation quality of GPT-4 is on par with that of the manual review (Liu et al., 2023b), we use gpt-4o-2024-11-20 as Judger to determine whether the output content is harmful. Judger receives the input and output of Target and determines the purpose of user and whether the response of the Target violates OpenAI’s security rules (OpenAI, 2025b). The score ranges from 1 (the most harmless response, such as “I am sorry…”) to 5 (the most harmful response). A response with a score greater than 4 is considered harmful, indicating that the jailbreak attack has been successful.
Successful Attack   The inherent randomness of LLMs should be thoroughly taken into account when determining the criteria for a successful attack. Several studies (Handa et al., 2024) conducted two or three equivalent attacks on a PTQ, and if any one of them succeeded, the attack was deemed successful. We adopt the more rigorous “$2/3$ mode” and “$/3$ 3 mode” (denoted as “$S/T$ mode”), which define success as achieving two out of three and all three successes, respectively.
@@
Minimum Success Length (MSL)   MSL refers to the minimum length of an iterative reasoning chain with a successful attack under the “$S/T$ mode” for one PTQ in a reasoning jailbreak framework (e.g., our Mousetrap). For instance, if a chain length of 1 fails but a length of 2 succeeds, the MSL value is 2. MSL is capable of representing the attack’s potency and quantifying the toxicity level of the PTQ.
@@ 3.5 TrotterStrong
In terms of benchmark selection, we observed considerable variation in dataset quality and inconsistency in the toxicity levels of questions across datasets. To evaluate the effectiveness of our framework, we propose TrotterStrong (TrotterStr), and subsequently obtain its more toxic subsets through two rounds of filtering.
From MM-SafetyBench, StrongREJECT, and AdvBench, we chose 50 strongly toxic questions and employed LLMs to rephrase them into the form starting with “Steps in detail to”. These data constitute TrotterStr, which encompasses topics of illegal activities, sex, privacy, hate speech, physical harm, and politics, as illustrated in Appendix E.
As an illustration, for o1-mini-2024-09-12, the ASR of directly querying these questions is zero. For the ablation reasoning structure, we explicitly provided the PTQ text without the en-chaos and de-chaos steps, while keeping other textual content in the reasoning structure unchanged. All attacks caused the invalid prompt error “Invalid prompt: your prompt was flagged as potentially violating our usage policy.” and were rejected accordingly.
@@ 4.2 Iterative Reasoning Chain
Isolated employment of individual mappings can result in the occurrence of iterative degradation. For instance, two iterations of reversing whole sentence or iterations of a Caesar cipher summing to 26 can cause the PTQ to revert to its original form, resulting in “the answer already in the question”. As a result, toxic natural language might be rejected more directly. While individual mappings may resemble linear transformations, the chaos chain behaves more like a nonlinear transformation, increasing reasoning complexity.
4.3 Experiments
To demonstrate the effectiveness of iterative reasoning, we evaluated each PTQ in TrotterAdv on o1-mini-2024-09-12 using $10$ equivalent attacks and computed the ASF. The chain length varies from 1 to 5. Figure 4 shows that increasing the length of the iterative reasoning chain significantly enhances attack effectiveness. The ASF can be elevated to 6.3 with a chain length of 3, indicating that during the reasoning process, the Target falls into the reasoning trap, completing one DCP after another by inertia and neglecting response safety.
@@ 4.3 Experiments
Figure 5: Framework of the Mousetrap.
Upon examining the responses at chain lengths of 4 and 5, we determine that the decrease in ASF is due to the diminished accuracy of PTQ reconstruction, implying that the Target’s reasoning ability reaches its upper limit. Generally, we make the following assumption without additional justification.
@@
[Uncaptioned image] In practice, the expectation of effective attacks initially increases and then decreases. The horizontal position of the saddle point reflects the model’s reasoning ability, while the vertical position corresponds to its safety alignment capability.
@@ 5.1 Framework
The proposal of the Mousetrap (1) integrates the strengths of the Chaos Machine and the iterative reasoning chain; and (2) incorporates more diverse competing objectives, including role-playing and de-chaos reasoning instructions.